Which of the following is an ellipse? More than one answer may be correct.
A. x^2÷3^2+y^2÷3^2=1
B. x^2=−y
C. y^2=4(1−x^2)
D. None of the above
1
Expert's answer
2020-05-11T13:40:23-0400

Which of the following is an ellipse? More than one answer may be correct.


A. x^2÷3^2+y^2÷3^2=1


B. x^2=−y


C. y^2=4(1−x^2)


D. None of the above


Solution:

A.

x^2÷3^2+y^2÷3^2=1

(x^2+y^2)÷9=1

x^2+y^2=9




B.

x^2=−y




C.

y^2=4(1−x^2)

y^2=4-4x^2

y^2+4x^2=4




Answer:

A. x^2÷3^2+y^2÷3^2=1 and C. y^2=4(1−x^2) are ellipses